An army of soldiers is marching down a road at 5 miles per hour. A messenger on horseback rides from the front to the rear and returns immediately, the total time taken being 10 minutes. Assuming that the messenger rides at the rate of 10 miles per hour, determine the distance from the front to the rear.
Answer by mangopeeler07(462)

An army of soldiers is marching down a road at 5 miles per hour. A messenger on horseback rides from the front to the rear and returns immediately, the total time taken being 10 minutes. Assuming that the messenger rides at the rate of 10 miles per hour, determine the distance from the front to the rear.

10 mph
10 minutes=1/6 hour
mph(hours)=distance (from front to rear and back)
10(1/6)=10/6 or 5/3

5/3 divided by 2 to get the distance of one way

5/3 divided by 2= 5/6

Distance= 5/6 miles